-A pole that is 2.6m tall casts a shadow that is 1.22 long. At the same time, a nearby tower casts a shadow that is 37.25m long. How tall is the tower? Round your answer to the nearest meter
Answer by Cromlix(4381)

Similar triangle problem.
Considering the length of the pole and the height of the tower
as one ratio and the length of the pole's shadow and the tower's shadow as the other ratio:-
2.6m/ x = 1.22m/37.25m
Cross multiply
1.22 * x = 2.6 * 37.25
x = (2.6 * 37.25)/ 1.22
Height of tower = 79 metres.
Hope this helps
:-)